# V2X-R: Cooperative LiDAR-4D Radar Fusion with Denoising Diffusion for 3D Object Detection

### v2

@@ 5.1 Experimental Details and Metrics
We used 8,084/829/3,166 frames for training/ validation/ testing in our V2X-R dataset, ensuring there is no overlap in the intersection of the training/validation/testing sets. For each frame, we ensure that the minimum and maximum numbers of CAVs are 2 and 7 respectively. We use Adam optimizer with lr = 1e-3, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999. Following [54], we select a vehicle as the ego vehicle for evaluation. Detection performance is evaluated near the ego vehicle in a range of x $\in$ [0, 140]m, y $\in$ [-40, 40]m. We set the broadcast range among CAVs to be 70 meters. In addition, since the 4D radar sensor only provides front-view data, all of our evaluation results are in the camera FOV of the ego vehicle.
@@ 5.2 Benchmark Models
Where2comm [12] NeurIPS2023 85.2/83.1/65.9 91.6/88.5/80.5
@@
Where2comm [12] NeurIPS2023 71.6/67.2/42.9 80.4/77.3/56.7
@@
1.Where2comm [12] L 61.39 60.62 52.11 72.63 70.19 52.60 86.25 84.99 73.78
2.Where2comm [12] L+4DR 85.93 82.19 65.27 85.17 81.45 63.91 91.40 90.26 80.90
3.Where2comm w/MDD L+4DR 87.71 85.82 69.43 87.11 84.30 66.67 91.42 90.34 80.58
Improvement (2-1) +24.54 +21.57 +13.16 +12.54 +11.26 +11.31 +5.15 +5.27 +7.12
Improvement (3-2) +1.78 +3.63 +4.16 +2.04 +3.01 +2.97 +0.02 +0.08 -0.32
@@ 5.3 Benchmark Analysis
Cooperative 4D radar-based 3D object detectors. We also pioneered the exploration of cooperative 4D radar-based methods on the V2X dataset. As shown in Table 3, the performance of 4D radar-based models is generally lower than that of LiDAR-based models in Table 2. This can be attributed to the significantly lower resolution of 4D radar than LiDAR, which is a limitation hindering the independent use of 4D radar in single-agent scenarios. However, we also observed that cooperative 4D radar-based methods can achieve the excellent performance of 85.8% 3D mAP@testing (IoU=0.3) (CoBEVT [55]). This shows that 4D radar can achieve independent perception through cooperative communication, proving the research significance of the cooperative 4D radar-based approach.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
Performance comparison under different simulated weather on V2X-R dataset. We selected three models, AttFuse [54], L4DR [16], and Where2comm [12], which include SA2MA and SM2MM fusion strategies, to evaluate the performance under different weather conditions. As shown in Table 5, the LiDAR-4D radar fusion can achieve superior performance in various weather conditions. Where2comm and AttFuse achieved 24.54% and 17.17% 3D mAP@snow (IoU=0.3) improvement when compared to LiDAR-only (Improvement (2-1)), respectively. Moreover, our designed MDD module improved the basic LiDAR-4D radar fusion model performance by up to 5.73% (AttFuse, 3D mAP@fog) and 6.70% (L4DR, 3D mAP@snow) under adverse weather conditions. There experimental results fully validate the effectiveness of our MDD module. At the same time, the addition of our MDD module hardly affects the lower impact of normal weather, highlighting its reliability.
@@ 6 Conclusion and Discussion
Limitations and future work. While our work provides a strong foundation, deeper exploration of cooperative LiDAR and 4D radar fusion remains underdeveloped. A compelling research direction is fully utilizing multi-agent and multi-modal information for more robust 3D object detection. Another promising direction is to develop cooperative 4D radar-based or radar-camera perception systems.
@@ 7.3 Information of data collection
PFA-Net [50] 30 4 7 0.001 Multistep
RTNH [29] 15 4 7 0.001 Multistep
@@
AttFuse [54] 30 4 7 0.002 Multistep
Where2comm [12] 50 1 5 0.0002 Cosineannealwarm
SCOPE [60] 30 2 5 0.002 Multistep
@@
CoAlign [26] 20 2 7 0.002 Multistep
AdaFusion [17] 30 2 5 0.0005 Multistep
@@
InterFusion [42] 20 1 7 0.002 Multistep
L4DR [16] 30 2 7 0.002 Multistep

### v3

@@ 5.1 Experimental Details and Metrics
We used 8,084/829/3,166 frames for training/ validation/ testing in our V2X-R dataset, ensuring there is no overlap in the intersection of the training/validation/testing sets. For each frame, we ensure that the minimum and maximum numbers of agents are 2 and 5 respectively. We use Adam optimizer with lr = 1e-3, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999. Following [54], we select a vehicle as the ego vehicle for evaluation. Detection performance is evaluated near the ego vehicle in a range of x $\in$ [0, 140]m, y $\in$ [-40, 40]m. We set the broadcast range among CAVs to be 70 meters. In addition, since the 4D radar sensor only provides front-view data, all of our evaluation results are in the camera FOV of the ego vehicle.
@@ 5.2 Benchmark Models
Where2comm [12] NeurIPS2023 83.7/80.2/56.8 88.1/86.1/69.2
@@
Where2comm [12] NeurIPS2023 69.4/62.1/26.6 77.8/72.9/41.5
@@ 5.3 Benchmark Analysis
Cooperative 4D radar-based 3D object detectors. We also pioneered the exploration of cooperative 4D radar-based methods on the V2X dataset. As shown in Table 3, the performance of 4D radar-based models is generally lower than that of LiDAR-based models in Table 2. This can be attributed to the significantly lower resolution of 4D radar than LiDAR, which is a limitation hindering the independent use of 4D radar in single-agent scenarios. However, we also observed that cooperative 4D radar-based methods can achieve the excellent performance of 86.4% 3D mAP@testing (IoU=0.3) (CoBEVT [55]). This shows that 4D radar can achieve independent perception through cooperative communication, proving the research significance of the cooperative 4D radar-based approach.
@@ 5.4 Multi-modal Diffusion Denoising Analysis
Performance comparison under different simulated weather on V2X-R dataset. We selected three models, AttFuse [54] and L4DR [16], which include SA2MA and SM2MM fusion strategies, to evaluate the performance under different weather conditions. As shown in Table 5, the LiDAR-4D radar fusion can achieve superior performance in various weather conditions. AttFuse achieved 17.17% 3D mAP@snow (IoU=0.3) improvement when compared to LiDAR-only (Improvement (2-1)), respectively. Moreover, our designed MDD module improved the basic LiDAR-4D radar fusion model performance by up to 5.73% (AttFuse, 3D mAP@fog) and 6.70% (L4DR, 3D mAP@snow) under adverse weather conditions. Meanwhile, our MDD module hardly affects the normal weather performance, which validates the effectiveness and reliability of our MDD module.
@@ 6 Conclusion and Discussion
Limitations and future work. While our work provides a strong foundation, deeper exploration of cooperative LiDAR and 4D radar fusion remains underdeveloped. A compelling research direction is the full utilization of multi-agent and multi-modal information for more robust 3D object detection.
@@ 7.3 Information of data collection
PFA-Net [50] 30 4 5 0.001 Multistep
RTNH [29] 15 4 5 0.001 Multistep
@@
AttFuse [54] 30 4 5 0.002 Multistep
Where2comm [12] 15 1 5 0.0002 Cosineannealwarm
SCOPE [60] 15 2 5 0.002 Multistep
@@
CoAlign [26] 20 2 5 0.002 Multistep
AdaFusion [17] 15 2 5 0.0005 Multistep
@@
InterFusion [42] 20 1 5 0.002 Multistep
L4DR [16] 30 2 5 0.002 Multistep